OpenStudy (anonymous):

How do i parametrize this? x^2 + y^2 + z^2 = 4(x+y) x+y=4

OpenStudy (anonymous):

allow x= t solve your bottom and sub i'm assuming

OpenStudy (anonymous):

i did horrible on this test so ... lol

OpenStudy (anonymous):

\[x ^{2}+ y ^{2} + z ^{2} = 4(x+y) \rightarrow (x-2)^{2} +(y-2)^{2}+z^{2} = 8 \] Then i do x = t in the second equation and get y = 4-t, replacing in the fisrt... do you think thats correct? z ends a bit ugly

OpenStudy (anonymous):

not sure how you got x-2?

OpenStudy (anonymous):

parametrize wont alter the equation i believe and you just made your center offset

OpenStudy (anonymous):

is x+y=4 you parameter?

OpenStudy (anonymous):

i was saying \[x^2+y^2+z^2=4(4)=16\] let t=x y=4-x y=4-t

OpenStudy (anonymous):

\[z^2=16-t^2-(4-t)^2\]

OpenStudy (anonymous):

\[z^2=16-16-t^2-t^2+8t=-2t^2+8t\]

OpenStudy (anonymous):

\[z=\sqrt{-2t^2+8t}\]

OpenStudy (anonymous):

so parameter would be \[t^2,(t-4)^2,\sqrt{-2t^2+8t}\]

OpenStudy (anonymous):

now i'm like around 50 on this one cause i failed this test lol

OpenStudy (anonymous):

Exactly that is what i ended with, but that square root seemed a little ugly to me, but now that you confirm it, im more confident, however, i think i can get some other things from this... let me check... i think there are 2 curves, one with -square root and the other with +squereroot....

OpenStudy (anonymous):

well the parameter they picked was ugly to begin with... if you were to pick it you could find a much more simple one

OpenStudy (anonymous):

both positive and negative i believe you have what looks like a plane intersecting a sphere

OpenStudy (anonymous):

or line

OpenStudy (anonymous):

Well i think i got it, the final answer would be: \[\Gamma _{1}: x= t // y = 4-t // z = \sqrt{-2t^{2}+8t}\] 0 <= t <= 4 \[\Gamma _{2}: x= t // y = 4-t // z = -\sqrt{-2t^{2}+8t}\] 0 <= t <= 4 The only thing that changes is the minus sign in t.

OpenStudy (anonymous):

in z i mean

OpenStudy (anonymous):

Thanks for your help.